OpenStudy (anonymous):

Find the solution of the linear equation. 2/3(6 - 3x) = (1 - x) + (3 - x) A. x = -8 B. x = 8 C. infinitely many solutions D. no solution

OpenStudy (anonymous):

Find the solution of the linear equation. 4a - 11 = 2a + 2(a + 6) A. a = 1 B. a = 12/12 C. infinitely many solutions D. no solution